# Adversarial Markov Games: On Adaptive Decision-Based Attacks and Defenses

### v1

@@ VI Discussion
With AI-enabled decision-making becoming more pervasive and accessible, in domains like governance, finance, employment, and of course cybersecurity, more and more decisions are delegated to AI which becomes increasingly accountable for upholding safety and ethical constraints. As trustworthy AI is vital for the healthy functioning of whole ecosystems, this study highlights several security risks but also potential mitigations in these inherently black-box environments.
@@
Our work has several implications for performing robust inference in the wild. While adversarial training remains the most reliable defense, the amount of robustness it imparts will vary and even be insufficient. We demonstrated how AI-enabled systems are susceptible to adaptive adversaries that devise new evasive techniques and control them jointly with other attack parameters. This has been achieved in the fully black-box case and against active defenses. Even more concerning is the level of threat that such systems face from AI-enabled adversaries, as it is straightforward to generalize Theorem III.1 to any other domain or modality. This rekindles the proverbial arms race, where as a consequence defenses should also be backed by equally capable AI.
Finally, our study highlights some aspects of the nature of adversarial examples. In black-box interactive environments they can be generalized by the notion of adversarial behavior, something that is more evident in domains where visual affinity is irrelevant, e.g., malware detection. Adversarial examples (and behavior) bring essence into question, what makes something, something; thus they can be approached from various epistemic perspectives. Computational approaches concern themselves with how to effectively acquire them; while their inherently competitive nature is already reflected in seminal works like shortcut learning [23] and invariant risk minimization [1]. Even in phenomenology – the philosophical study of the structures of experience and consciousness – the eidetic variation technique involves imagining and varying the features of an object under investigation to discover the (in)essential ones [31]. What these disparate approaches bring to light is that towards more general and robust AI researchers will have to reflect on the causal structure of perceptions and preconceptions, as well as on ways to automate this process.
Limitations. We note that while adversarial training has limitations, the robustness it provides is against all kinds of adversarial examples (under the same $\ell_{p}$-norm), while active defenses protect primarily against attacks they were trained with. As complementary approaches however, they can be used jointly. To keep the amount of evaluations practical, we narrowed the scope to targeted attacks and to $\ell_{2}$ as the more suitable norm for visual similarity. Targeted attacks are strictly more difficult to perform than untargeted, while in binary classification targeted and untargeted coincide; the latter is also the prevalent mode in cybersecurity contexts. Our framework, however, can accommodate any adversarial goal or metric.
Another simplifying assumption we make is that only one attack can take place at a time; however, the queuing methodology we use for incoming queries is readily extensible to handle concurrent attacks. Other possibilities to consider is how well an active defense handles attack methodologies it was not trained on, and to extend its training to multiple ones.
Finally, in the evaluation we focus on the breadth of scenarios and environments where agents learn concurrently, thus the number of datasets we experiment with is limited. However, we back our empirical study with an extensive theoretical analysis that gives support to our findings generalizing to any dataset or application domain.
@@
Future Work. The AMG framework we introduce is general by design and can accommodate the learning of optimal offensive and defensive policies in any domain of interest beyond image classification. A promising path for future research is the extension of adaptive attacks and defenses to other domains and modalities, e.g. tasks like malware, bot, and network intrusion detection, precisely because our approach circumvents the obstacle of computing and mapping gradients to feasible perturbations and operates instead directly on the problem space [39]. Another compelling but formidable challenge is automating the adaptive evaluations in AML, that is adapting beyond a specification by inventing instruments to bypass defenses and thus imparting controllability to adversarial tasks. Finally, in our work we considered opponent agency as part of the environment; other domains, like malware analysis, might benefit from explicit opponent modeling.
@@ VII Conclusion
With decision-based attacks becoming more pervasive in multiple domains, every AI-based decision-making process that exposes a queryable interface is inherently vulnerable. To aggravate matters, this vulnerability cannot be mitigated by employing model hardening approaches like adversarial training alone. To fully defend in the presence of such attacks, active and adaptive defenses are necessary, and we demonstrate how optimal defensive policies can be learned. However, the existence of such defenses elicits in turn adaptive attacks which are able to recover part of their original performance.
We perform a theoretical and empirical investigation of decision-based attacks and defenses under a unified framework we name “Adversarial Markov Games” (AMG). In self-adaptive, we introduce a novel twofold definition of adaptive: both devising new methods of outmaneuvering opponents and adapting one’s operating policy with respect to other agency in the environment. Furthermore, our adversarial policy gradient theorem indicates that any combination of adversarial goals, be it performance, stealthiness, disruption, can be optimized in a gradient based manner, even in the complete black-box case and in any modality. As new attacks and defenses appear (and get broken) regularly, our evaluation methodology is applicable to any within the outlined scope by transforming the existing arms-race through their self-adaptive versions, thus ensuring accurate and robust assessment of their performance.
The AMG framework we introduce helps us reason on and assess the vulnerabilities of AI-based systems by disentangling the inherently complex and non-stationary task of learning in the presence of competing agency. We can model it as part of the environment, and thus simplify the task by computing a best response against the observed behavior. This is an important outcome for cybersecurity domains: as long as proper threat modeling is carried out, one can readily employ RL algorithms in order to devise optimal defenses; but only after they devised optimal attacks too. w
@@ Appendix A Proofs
For a more intuitive understanding of the proofs, we first provide a high-level description of the attack fundamentals. BAGS [12] performs a random walk along the boundary between the adversarial and the non-adversarial regions, by first taking a random step orthogonal to the original image direction, then a source step towards it. The randomness in the directions searched is reduced by utilizing Perlin noise and masks computed on the difference between starting and original samples. HSJA [15] operates in 3 stages: a binary search that places the current best adversarial on the decision boundary, an estimation step that computes the gradient at that point of the boundary, and projection step along the estimated gradient. These steps repeat until convergence.
@@
Recall that we assumed $C(x)=\operatorname*{arg\,max}F_{c}(x)$, then the updates in Eq. (16) are guaranteed by Theorem 1 of HSJA [15] to converge to a stationary point $x_{b}$ of Eq. (5). Given a standard threshold $\epsilon$ on imperceptibility and over $N$ adversarial episodes, this implies that $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]<\epsilon$. As $C$ is the only term in the expression (16) that is affected by the model, we reach a contradiction and that an alternative classifier $C^{\prime}$ is required, with $C(x)=D(F_{c}(x))$. As the discriminant function $F_{c}$ cannot change without retraining, it follows that $D^{\prime}\neq\operatorname*{arg\,max}$, so that for the adversarial example $x_{t}$ misclassified as $c_{0}$, Eq. 4 can return -1:
@@
[Theorem III.1]
Proof.
Let $H$ be a distance-based detector that based on a stateful representation $\mathcal{S}(\tau,x_{t})$ decides if a query $x_{t}$ is adversarial or not, and we assume the candidate generation policy $\pi_{\theta}^{\mathcal{A}}$ can influence this representation through intrinsic parameters like source step size, or applying transformations the underlying model is invariant to (V). For a predefined threshold $\delta$, if $\mathcal{S}(\tau,x_{t})<\delta$ then $H(x_{t})=1$ indicating that $x_{t}$ is part of an adversarial attack, otherwise $H(x_{t})=0$. Now for a number $N$ of attack episodes of fixed length $L$ and a reward function $r(\tau)=\sum_{t=1}^{L}(1-H(x_{t}))$, the objective to be maximized is the expected reward of the adversarial policy:
$\mathcal{J}(\theta)=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau)]=\frac{1}{N}\sum_{i=1}^{N}r(\tau)=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{L}(1-H(x_{t}))$ (18)
As it is not possible to compute the derivative of an expectation, by using the Policy Gradient Theorem [48] we can transform it to the expectation of the product of the reward and the gradient of the logarithm of the policy:
$\nabla\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau)]$ $=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau)\nabla\log\pi_{\theta}(\tau)]$ (19)
$=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[\sum_{t=1}^{L}(1-H(x_{t}))\nabla\log\pi_{\theta}^{\mathcal{A}}(\tau)]$
$=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{L}(1-H(x_{t}))\nabla\log\pi_{\theta}^{\mathcal{A}}(\tau)$
To maximize this expectation empirically we sample $e$ episodes from the environment, compute the gradient of the policy $\nabla\log\pi_{\theta}^{\mathcal{A}}(\tau)$ and then update the policy with a learning step $a$, $\theta^{\prime}=\theta+\alpha\cdot\sum_{0}^{e}r(\tau)\nabla\log\pi_{\theta}^{\mathcal{A}}(\tau)$. After this process has converged through gradient ascent, and as the maximum in Eq. 18 is attained for $\sum_{i=1}^{N}\sum_{t=1}^{L}H(x_{t})=0$, the obtained policy $\pi_{\theta}^{\mathcal{A}}$ will correspond to the optimally evasive one $\mathcal{E}$. ∎
@@ Appendix B On States & Rewards
[States]. In POMDPs a single observation cannot constitute a Markovian state as typically there is some of dependence on the state history. Partial observability is handled by either using recurrent architectures for the policy and value networks, or by engineering a state that includes past information: in this work we opt for the later.
For BAGS, the adversary uses an 8-dimensional state representation with the following information normalized in the range $[0,1]$ : current amount of queries $i$, average queries that are adversarial $a$, the initial gap $g$, the current gap $d$, the location $l=\frac{d}{g}$, the slope $s=m-l$ where $m$ is a moving average of the location, the frequency of improvement $f$, and $r$ which is a moving average of the perturbation reduction $n$, In HSJA the state representation is slightly different: $r=\frac{n}{g}$, and $f=\frac{1}{j}$ with $j$ number of jump steps in last iteration.
Regarding the defense, for HSJA (and to a lesser extent BAGS) it has been difficult to engineer a representative state for policies to effectively learn on. The knowledge of the attack internals and fundamentals, geometric properties and distances, model activations and logits, and any combination thereof, did not suffice. Ultimately we turned to pure computation to learn a state representation for the defensive agent. This representation is a 64-dimensional embedding of a CNN trained with triplet loss, on data generated by HSJA and benign queries, where the input is a tensor where the last query is subtracted from the 25 most recent adversarial queries and then stacked together.
[Rewards].
The concrete definitions of the rewards for each type of agent are:
@@
Appendix C Models & Hyperparameters
The image classification models we use are ResNet-20 for CIFAR-10 and a standard 2 convolutional / 2 fully-connected layer NN for MNIST. For adversarially training models, we follow the canonical approach as described in [53]: the model is trained for 20 epochs, where the first 10 are trained normally and the last 10 on batches containing additional adversarial examples generated with 40 steps of PGD. For learning the similarity space, that is the metric space where defensive agents control the radius of interception around which a query is adversarial or not, we use a Siamese CNN. This network is trained with contrastive loss, where dissimilar examples are generated by adding Gaussian noise and performing evasive transformations on the input from the list in Table V. For the PPO agents trained for each scenario, we use the open source library Stable Baselines 3  [^2] https://github.com/DLR-RM/stable-baselines3. Policies are parameterized by a two fully-connected layer NN; the hyperparameter search space is shown in Table VI.
TABLE VI: Hyperparameter space for the PPO agent.

### v2

@@ VI Discussion
Our work has several implications for performing robust inference in the real-world. While adversarial training remains the most reliable defense, the empirical robustness it imparts will vary and even be insufficient. We note that this robustness is against all adversarial examples under the same $L_{p}$-norm; active defenses protect only against querying attacks, but as they do transfer between attacks (cf. Scen. 12) they can be used jointly as complementary approaches. We demonstrated how AI-enabled systems are susceptible to adaptive adversaries that devise new evasive techniques and control them jointly with other attack parameters. This has been achieved in the fully black-box case and against active defenses. Notably, the level of threat that adaptive adversaries pose against such systems is considerable, as it is straightforward to generalize Proposition III.3 to any other domain or modality. This rekindles the proverbial arms race, where as a consequence defenses should also be equally capable and adaptive.
Limitations. To keep the amount of evaluations practical, we narrowed the scope to targeted attacks and to $L_{2}$ as the more suitable norm for visual similarity. Targeted attacks are strictly more difficult to perform than untargeted, while for binary classification targeted and untargeted coincide; our framework, however, can accommodate any adversarial goal or metric. Another simplifying assumption we make is that only one attack can take place at a time; however, the queuing technique we use for incoming queries is readily extensible to handle concurrent attacks. While we demonstrate that our active defense does transfer between attacks, another possibility to explore is training the defense on multiple kinds of attacks. Finally, in our evaluation we focus on a wide range of adaptive and non-adaptive scenarios where agents learn and adapt interactively, thus limiting the number of datasets we experiment with; we back our empirical study however with an extensive theoretical analysis that supports the generality of our findings independent of context.
@@
Future Work. The AMG framework we introduce is general by design and can accommodate the learning of optimal offensive and defensive policies in any domain of interest beyond image classification. A promising path for future research is the extension of our adaptive attacks and defenses to other domains and modalities, for instance malware, bot, and network intrusion detection. This is specifically because our approach circumvents the main obstacle of mapping gradient-based perturbations to feasible objects (eg. binaries) and instead can function directly in the problem space [36]. Another compelling and formidable challenge is automating the adaptive evaluations in AML, that is adapting beyond a specification by inventing instruments to bypass defenses and thus imparting controllability to adversarial tasks. Finally, in our work we considered opponent agency as part of the environment; other domains, like malware detection, might benefit from explicit opponent modeling.
@@ VII Conclusion
With adaptive, decision-based attacks becoming more pervasive in multiple domains, every AI-based decision-making process that exposes a queryable interface is inherently vulnerable. To aggravate matters, this vulnerability cannot be mitigated by employing model hardening approaches like adversarial training alone. To fully defend in the presence of such attacks, active and adaptive defenses are necessary, and we demonstrate how optimal defensive policies can be learned. However, the existence of such defenses elicits in turn adaptive attacks which are able to recover part of their original performance.
We perform a theoretical and empirical investigation of decision-based attacks and stateful defenses under a unified framework we name “Adversarial Markov Games” (AMG). In self-adaptive, we introduce a novel twofold definition of adaptive: both devising new methods of outmaneuvering opponents and adapting one’s operating policy with respect to other agency in the environment. Furthermore, our theoretical analysis highlights that any combination of adversarial goals, be it performance, stealthiness [52], or disruption, can be optimized in a gradient based manner, even in the complete black-box case and in any modality. As new attacks and defenses appear and get broken regularly, our evaluation methodology is generally applicable within the outlined scope by turning any approaches in the existing arms-race self-adaptive and thus ensuring accurate and robust assessment of their performance.
The AMG framework we introduce helps us reason on and properly assess the vulnerabilities of AI-based systems, disentangling the inherently complex and non-stationary task of learning in the presence of competing agency. By modeling the latter as part of the environment, we can simplify this task by computing a best response to the observed behavior. This is a significant implication for the security of AI-based systems: as long as proper threat modeling is carried out, one can readily employ RL algorithms in order to devise optimal defenses, but only after they devised optimal attacks too.
@@ Appendix A Proofs
For a more intuitive understanding of the proofs, we provide a high-level description of the attack fundamentals. BAGS [29] performs a random walk along the boundary between the adversarial and the non-adversarial regions, by first taking a random step orthogonal to the original image direction, then a source step towards it. The randomness in the directions searched is reduced by utilizing Perlin noise and masks computed on the difference between starting and original samples. HSJA [7] operates in 3 stages: a binary search that places the current best adversarial on the decision boundary, an estimation step that computes the gradient at that point of the boundary, and projection step along the estimated gradient. These steps repeat until convergence.
@@
In Eq. (16), the iterates $x_{t}$ are guaranteed by Theorem 1 of HSJA [7] to converge to a stationary point $x_{b}$ of Eq. (5), that is $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]<\epsilon$, for $\epsilon$ a standard imperceptibility threshold and $N$ adversarial episodes. We thus reach a contradiction, and as in Eq. (16) the only model-dependent term is $C(\cdot)$, an alternative classifier $C^{\prime}$ is required. With $C(x)=D(F_{c}(x))$, and the discriminant function $F_{c}$ unable to change without retraining, it follows that $D^{\prime}\neq\operatorname*{arg\,max}$, so for adv. example $x_{t}$ misclassified as $c_{0}$, Eq. 4 can return -1:
@@
[Proposition III.3]
Proof.
Let $\pi_{\theta}^{\mathcal{A}}$ be the adversarial policy generating queries $x_{t}$, in $N$ episodes $\tau_{i}$ of length $L$. The defense $\pi_{\phi}^{\mathcal{D}}$ (7), upon receiving a query $x_{t}$ outputs a decision $\alpha_{t}=\pi_{\phi}^{\mathcal{D}}(x_{t},s_{t}^{\mathcal{D}})\in\{0,1\}$, with $1$ and 00 indicating rejection and acceptance respectively. The goal of the adversary is to find the parameters $\theta^{*}$ that maximize the expected reward:
$\mathcal{J}(\theta)=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau_{i})]=\frac{1}{N}\sum_{i=1}^{N}r(\tau_{i})=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{L}(1-\alpha_{t})$ (21)
The gradient of $\mathcal{J}(\theta)$ with respect to the policy parameters $\theta$ is the direction of steepest ascent for Eq. (21). Through the Policy Gradient Theorem [53] we can express the gradient of the expected reward in terms of the gradient of the log-likelihood of the policy, weighted by the reward:
$\nabla_{\theta}\mathcal{J}(\theta)$ $=\nabla_{\theta}\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau_{i})]$ (22)
$=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}\left[r(\tau_{i})\nabla_{\theta}\log\pi_{\theta}(\tau_{i})\right]$
$=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}\left[\sum_{t=1}^{L}(1-a_{t})\nabla_{\theta}\log\pi_{\theta}^{\mathcal{A}}(\tau_{i})\right]$
$\approx\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{L}(1-a_{t})\nabla_{\theta}\log\pi_{\theta}^{\mathcal{A}}(\tau_{i})$
The gradient is thus estimated by sampling $N$ episodes from the policy $\pi_{\theta}^{\mathcal{A}}$ to compute Eq. (LABEL:eqn:rew_app). To maximize the expectation, we iteratively update the policy parameters $\theta$ using gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}(\theta)$. With $\eta>0$ the learning rate, this process converges to $\theta^{*}$. Recall that Eq. (21) attains its maximum for $\sum_{i=1}^{N}\sum_{t=1}^{L}a_{t}=0$, therefore the converged policy $\pi^{\mathcal{A}}_{\theta^{*}}$ will correspond to the optimal evasive policy.
@@ Appendix B On States & Rewards
[States]. By merit of operating in black-box environments, the optimization process for both offensive and defensive agents can be described as a POMDP. To handle this partial observability, the states we engineer incorporate past information. For BAGS, the adversary uses an 8-dimensional state representation with the following information normalized in the range $[0,1]$ : current amount of queries $i$, average queries that are adversarial $a$, the initial gap $g$, the current gap $d$, the location $l=\frac{d}{g}$, the slope $s=m-l$ where $m$ is a moving average of the location, the frequency of improvement $f$, and $r$ which is a moving average of the perturbation reduction $n$, In HSJA the state representation is slightly different: $r=\frac{n}{g}$, and $f=\frac{1}{j}$ with $j$ number of jump steps in last iteration.
For the defense, in HSJA (and to a lesser extent BAGS) it has been difficult to engineer a state for policies to effectively learn on. The knowledge of the attack internals and fundamentals, geometric properties and distances, model activations and logits, and any combination thereof, did not suffice. Ultimately we decided to learn a representation instead. This representation is a 64-dimensional embedding of a CNN trained with triplet loss, on data generated by HSJA and benign queries, with the input being a tensor of the last query subtracted from the 25 most recent adversarial queries and then stacked.
[Rewards]. The concrete definitions of the rewards for each type of agent are:
@@
Appendix C Base Rate of Attacks
In all the evaluations so far we use a fixed probability $P(adv)=0.5$ that an incoming query is adversarial. To assess how our adaptive stateful defenses (Scenarios 4 & 6) perform in the complete absence of attacks, we evaluate them with $P(adv)=0$ without retraining; the results are shown in Table VI. We observe a small reduction in the accuracy on clean samples that can be attributed to the considerably different base rate of adversarial and benign queries. Note however that as the probability of adversarial queries is an intrinsic property of each environment, if the base rate of attacks changes the defensive agents can be retrained to adjust to it.
TABLE VI: Clean accuracy on CIFAR-10 for scenarios 4 & 6, where $P(adv)$ denotes the probability that a query is part of an attack.
Adv. Trained $P(adv)$ BAGS4 BAGS6 HSJA4 HSJA6
0.5 91.55 91.38 91.59 91.62
0.0 90.91 90.95 90.48 90.86
0.5 87.61 87.50 87.58 87.68
0.0 87.02 87.11 86.70 86.98
Appendix D Models & Hyperparameters
The image classification models we use are ResNet-20 for CIFAR-10 and a standard 2 convolutional / 2 fully-connected layer NN for MNIST. For adversarially training models, we follow the canonical approach as described in [5]: the model is trained for 20 epochs, where the first 10 are trained normally and the last 10 on batches containing additional adversarial examples generated with 40 steps of PGD. For learning the similarity space, that is the metric space where defensive agents control the radius of interception around which a query is adversarial or not, we use a Siamese CNN. This network is trained with contrastive loss, where dissimilar examples are generated by adding Gaussian noise and performing evasive transformations on the input from the list in Table V. For the PPO agents trained for each scenario, we use the open source library Stable-Baselines3  [^3] https://github.com/DLR-RM/stable-baselines3. Policies are parameterized by a two fully-connected layer NN; the hyperparameter search space is shown in Table VII.
Blacklight & OARS. For evaluating Blacklight [23] and OARS [24] we use their default hyperparameters as those are provided in the publicly available implementations. In particular, as OARS spends 200 extra queries per episode to adapt the proposal distribution, we add those on top of b the evaluation budget. Additionally, as our defense is not rejection based, we replace the rejection decision with a non-adversarial one.
TABLE VII: Hyperparameter space for the PPO agent.